OpenStudy (anonymous):

Y=3x+6 is? Choose one answer. a. What is the slope intercept form of the equation ? b. What is the slope intercept form of the equation ? c. What is the slope intercept form of the equation ? d. What is the slope intercept form of the equation ?

OpenStudy (anonymous):

just realize I didn't type the whole questions a is 3x-y=-6 b is x-3y=3 c is x+3y=-6 and d is 9x+3y=6

OpenStudy (anonymous):

a

OpenStudy (chriss):

a. y=3x-6 b. y=(1/3)x+1 c. y=(-1/3)x-2 d. y=-3x+2

OpenStudy (chriss):

so the answer is a